# MK-SGN: A Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation for Skeleton-based Action Recognition

### v4

@@ 3.4.2 Inner-layer feature distillation
The inner-layer features of heterogeneous GCN and SGN models differ significantly. While GCN typically processes continuous decimal values, SGN operates on sparse discrete spike sequences. To address this discrepancy, we propose the FTM, as illustrated in Figure 4(b). The FTM is an learnable module that transforms the continuous intermediate features of GCN into spike-form features, making them compatible with SGN.
The FTM concatenates multimodal features along the channel dimension and fuses them using depth-wise separable convolution. Subsequently, continuous features are expanded. Finally, spike-form features are generated through Conv2D, BN, and SN layers, ensuring compatibility between the teacher GCN and student SGN architectures while preserving the valuable structural and semantic information embedded in the GCN features.
In our distillation framework, the teacher GCN is composed of 10 GC-TC (Graph Convolution-Temporal Convolution) units, while the student SGN consists of 10 SGC-STC (Spiking Graph Convolution-Spiking Temporal Convolution) units. For feature distillation, we leverage intermediate-level features ($T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}$) and high-level features ($T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}$). The intermediate features are represented as: $T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}\in\mathbb{R}^{2C\times\frac{T}{2}\times V},$ while the high-level features are:$T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}\in\mathbb{R}^{4C\times\frac{T}{4}\times V}.$ These layers are selected because the expansion of channel dimensions and the changes in stride within the TCN network allow them to capture diverse temporal relationships and richer feature information. Taking the intermediate 5th layer as an example, the multimodal teacher features are first concatenated and fused as follows:
@@
To align the student SGN features with the teacher GCN features, we use cosine similarity as the loss function. For the intermediate 5th-layer spike-form feature and the corresponding SGN 3rd-layer spike-form feature ($S_{3}$), the loss is defined as:
$L_{\mathrm{fkd}_{1}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm5}}^{j}\cdot S_{3}^{j}}{\|T_{\mathrm{Sm5}}^{j}\|\cdot\|S_{3}^{j}\|}\right),$ (30)
where $J$ is the number of samples. Similarly, for the high-level 8th-layer spike-form feature and the SGN 5th-layer feature ($S_{5}$):
$L_{\mathrm{fkd}_{2}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm8}}^{j}\cdot S_{5}^{j}}{\|T_{\mathrm{Sm8}}^{j}\|\cdot\|S_{5}^{j}\|}\right).$ (31)
@@
$L_{\mathrm{fkd}}=\beta_{1}L_{\mathrm{fkd}_{1}}+\beta_{2}L_{\mathrm{fkd}_{2}},$ (32)
@@
Finally, the overall learning objective combines the inner-layer feature distillation loss with the task loss $L_{t}ask$ and the soft label distillation loss $L_{\mathrm{sdk}}$ as follows:
$L=\gamma_{1}L_{\mathrm{b}}+\gamma_{2}L_{\mathrm{sdk}}+\gamma_{3}L_{\mathrm{fkd}},$ (33)
@@ 4 Experiments
In this chapter, we present the implementation details, ablation studies, result comparisons, and visualizations, respectively. The details will be described in the following subsections.
@@ 4.1.1 Datasets
Cross-Subject Cross-View (CSet)
@@
NTU RGB+D 60 40,320 16,560 37,920 18,960 25
NTU RGB+D 120 120 63,026 53,370 54,471 61,925 25
NW-UCLA 10 - - 1,039 455 20
NTU-RGB+D is a large-scale dataset comprising 60 action classes and 56,000 video clips. It includes detailed skeleton data covering a wide range of daily activities, captured using a Kinect v2 sensor. The dataset provides 3D coordinates of 25 major body joints, enabling precise modeling of human movements. NTU-RGB+D is widely recognized as a standard benchmark in the action recognition field due to its variety of actions and large-scale data volume. The dataset is evaluated under two standard settings: Cross-Subject (CSub), where training and testing samples are divided based on the subjects performing the actions, and Cross-View (CView), where training and testing samples are divided based on the camera views.
NTU-RGB+D 120, an extended version of NTU-RGB+D, significantly expands the dataset to include 120 action classes and over 114,000 video clips. This dataset offers richer skeleton data, making it one of the largest and most comprehensive resources for skeleton-based action recognition. Its inclusion of a broader range of actions and more extensive data volume provides an ideal foundation for developing and evaluating advanced action recognition algorithms. Similarly, this dataset is evaluated under two standard settings: Cross-Subject (CSub) and Cross-Setup (CSet), where training and testing samples are divided based on the camera setups used during data collection.
NW-UCLA is a smaller dataset consisting of 10 action classes and 1,494 video clips captured from multiple viewpoints. It provides 3D coordinates of 20 major body joints, offering diverse skeleton data for assessing the robustness of action recognition models to viewpoint variations. The dataset poses a challenge for models to perform consistently across different perspectives. The evaluation is conducted under the Cross-View setting, where training and testing samples are divided based on the camera angles used for data capture.
@@ 4.1.2 Implementation Details
The experiments are conducted on a server equipped with an NVIDIA SMI V100 GPU with 30 GB of memory. The models are implemented using PyTorch[56] and SpikingJelly[55], which provide efficient frameworks for spiking neural networks. During training, we use two separate SGD optimizers: one dedicated to optimizing the SGN model and the other specifically designed to refine the FTM. This dual-optimizer strategy enhances the efficiency of the SGN and ensures effective feature transfer from the teacher to the student model.
@@
The training configurations for MK-SGN across different datasets are summarized in Table 3. We adopt an initial learning rate of 0.1, a weight decay rate of 1e-4, and set the batch size based on the dataset: 64 for NTU-RGB+D 60, 32 for NTU-RGB+D 120, and 128 for NW-UCLA. Training is conducted for 64 epochs with a step decay schedule, reducing the learning rate after 50 epochs.
@@
In the testing phase, only the SGN model is evaluated, isolating it from the dual-optimizer framework used during training. This ensures that the SGN’s performance reflects its learned capabilities without additional computational overhead.
@@ 4.2 Ablation Studies
To verify the effectiveness and efficiency of the proposed model, we first replace the activation layers of the classic one-stream 2S-AGCN model with LIF neurons. This modified model, called Base-SGN*, is then trained and tested for comparison. Additionally, we compare our proposed model with Spikformer, a classic SNN-based transformer architecture. For a fair comparison, we adjust the layer structure and hidden dimension of Spikformer to match the MK-SGN architecture. These adjustments ensure that the comparison focuses on the modeling approach while minimizing the influence of structural differences. Based on these baselines, we conduct ablation and comparison experiments to evaluate the performance of MK-SGN.
@@ 4.2.1 Theoretical Energy Consumption Calculation
In the context of SNNs, synaptic operations (SOPs) play a critical role in determining theoretical energy consumption. As one of the key components of our experimental evaluation, we focus on calculating SOPs and the energy consumption of both Base-SGN* and MK-SGN, highlighting the efficiency of the proposed model.
@@
Method Architecture Param Spike Time Step SMF Knowledge Distillation SOPs FLOPs Power Ntu-RGB+D 60
(M) (G) (G) ($mJ$) CSub(%) CView(%)
ANN (10-layer)ST-GCN 3.1 / / / / 3.48 16.01 81.5 88.3
@@
(6-layer)SpikformerICLR 2023 4.78 4 / / 1.69 24.07 2.17 73.9 80.6
@@
Following [27, 28, 29], the firing rate $r$ of each LIF neuron, is defined as the firing probability per spike time step. Using the firing rate $r$, the synaptic operations for a given block or layer $l$ are expressed as:
@@
where $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of multiply-and-accumulate(MAC) operations performed in $l$ block or layer. And $\operatorname{SOPs}$ are the number of spike-based accumulate (AC) operations.
@@
$\quad+E_{\text{AC}}\times\bigg{(}\frac{k\times(k-1)}{2}\mathrm{SOP}_{SMIC}+\sum_{n=n_{m}}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}$
@@
These formulations demonstrate how MK-SGN balances high performance with reduced theoretical energy consumption. SOPs, representing synaptic operations, serve as a critical metric for evaluating the energy efficiency of spiking-driven architectures.
@@ 4.2.2 The Effect of Different Components
To analyze the energy efficiency of the proposed MK-SGN and evaluate the contribution of each component within the model, we use theoretical formulas to estimate the number of operations(OPs) required for classifying a single action sample. Specifically, we calculate the FLOPs for ANN-based models, as well as the FLOPs for SNN on traditional GPUs and the theoretical SOPs on neuromorphic hardware. Using these operation counts, we further compute the energy consumption per action sample during the testing phase.
Additionally, we evaluate the classification accuracy of MK-SGN through comprehensive ablation studies conducted on the NTU-RGB+D 60 dataset. The results, including energy efficiency and accuracy comparisons, are summarized in Table 5.
@@
[…]
@@
Building on Base-SGN*, we compare its performance with the proposed MK-SGN. Even without knowledge distillation, the single-modality MK-SGN achieves significant accuracy improvements of 10.1% and 8.9% on XSub and XView settings, respectively, with only a minimal increase in OPs and theoretical power consumption. Compared to ST-GCN, although MK-SGN demonstrates a slight increase in FLOPs, this trade-off is offset by its extremely low SOPs value of 0.68, which highlights its suitability for deployment on neuromorphic hardware. Furthermore, MK-SGN reduces the parameter count by over 50% and decreases FLOPs and theoretical energy consumption by more than 70% compared to structurally similar models like Spikformer, while maintaining competitive computational efficiency.
@@
Furthermore, MK-SGN achieves an accuracy improvement of 0.6% on XSub and 0.3% on XView compared to Spikformer, demonstrating its superior balance of performance and efficiency. These results highlight the effectiveness of MK-SGN in achieving high accuracy with significantly reduced computational and energy requirements.
@@
In the case of late fusion in a multi-modal classification network, where $n_{n}$ represents the number of modalities, classifying a single action sample requires $n_{n}$ independent testing runs. This leads to a theoretical energy consumption scaled by $n_{n}$ times. Taking ST-GCN as an example, its theoretical OPs will become $4\times 3.48\,\text{G}$, and the Power will increase to $4\times 16.01\,\text{mJ}$.
To validate the effectiveness of the designed SMF module in avoiding the increased energy consumption caused by late fusion while simultaneously improving accuracy, we conduct more in-depth experiments, as shown in Table 6. Specifically, we incrementally integrate different modalities such as Joint (J), Joint+Bone (J+B), Joint+Bone+Joint-Motion (J+B+JM) and Joint+Bone+Joint-Motion +Bone-Motion(J+B+JM+BM)
[…]
These results validate that the SMF module effectively avoids the exponential increase in energy consumption caused by late fusion while consistently improving accuracy. Compared to single-modality fusion, the full multimodal fusion achieves an improvement of $1.6\%$ in accuracy for CSub, showcasing the robustness of the SMF module in efficiently integrating information across multiple modalities. This balance of efficiency and accuracy highlights the superiority of SMF over traditional late fusion methods.
@@
Soft Label Only - J 75.1
@@
We compare the accuracy of SGN models trained with and without teacher guidance to evaluate the impact of our proposed knowledge distillation method. Knowledge distillation consists of two components: soft label distillation (SKD) and inner-layer feature distillation (IFD). To further analyze their contributions, we conduct experiments with and without each component. Additionally, we investigate the impact of soft label distillation using different modality combinations, as well as the effect of inner-layer feature distillation applied at different levels (low-level, high-level, or both).
The results in Table 7 demonstrate the significant contributions of SKD and IFD to the accuracy improvements of SGN models. With the introduction of SKD, accuracy improves steadily as more modalities are incorporated. Single-modality SKD achieves 75.1%, while combining two and four modalities further enhances the performance to 75.4% and 75.6%, respectively. This progression highlights the importance of multimodal knowledge transfer from the teacher network to the student model.
For IFD, focusing on low-level features yields an accuracy of 74.9%, while high-level features achieve 75.3%, indicating that the more abstract information encoded in high-level features plays a slightly greater role in guiding the student model. However, combining both low-level and high-level feature distillation demonstrates their complementary effects, achieving a final accuracy of 75.9% when combined with SKD in the full knowledge distillation (Full KD) setup.
The results highlight the importance of leveraging both output-level and feature-level guidance. While SKD aligns the student’s outputs with the teacher’s multimodal predictions, IFD bridges the gap in intermediate representations, allowing the student to better emulate the teacher’s internal structure. The findings validate the proposed knowledge distillation framework as an effective means of improving SGN model performance through both comprehensive guidance and multimodal integration.
@@
We carefully evaluate the contributions of each component in the MK-SGN model, with results summarized in Table 5. The fully integrated MK-SGN achieves accuracies of 78.5% and 85.6% on XSub and XView settings, respectively, with a spike time step of 4. Compared to Spikformer, MK-SGN reduces the parameter count by 54.6% and FLOPs by 67.4%, significantly enhancing its feasibility for training on traditional GPUs. Moreover, MK-SGN decreases SOPs by 59.76% and theoretical energy consumption by 71.7%, while achieving accuracy improvements of 4.6% and 5.0%, demonstrating its ability to balance high performance with low energy consumption.

### v5

@@ 3.4.2 Inner-layer feature distillation
The inner-layer features of heterogeneous GCN and SGN models differ significantly. While GCN typically processes continuous decimal values, SGN operates on sparse discrete spike sequences. To address this discrepancy, we propose the FTM, as illustrated in Figure 4(b). The FTM is a learnable module that transforms the continuous intermediate features of GCN into spike-form features, making them compatible with SGN.
[…]
@@
To align the student SGN features with the teacher GCN features, we use cosine similarity as the loss function. For the converted spike-form feature from the multimodal GCN and the corresponding 3rd-layer spike-form feature in SGN ($T^{(3)}$), the loss is defined as:
$\mathcal{L}_{\mathrm{fkd}_{1}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm5}}^{j}\cdot{T^{(3)}}^{j}}{\|T_{\mathrm{Sm5}}^{j}\|\cdot\|{T^{(3)}}^{j}\|}\right),$ (29)
where $J$ is the number of samples. Similarly, for the high-level 8th-layer spike-form feature and the SGN 5th-layer feature ($T^{(5)}$):
$\mathcal{L}_{\mathrm{fkd}_{2}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm8}}^{j}\cdot{T^{(5)}}^{j}}{\|T_{\mathrm{Sm8}}^{j}\|\cdot\|{T^{(5)}}^{j}\|}\right).$ (30)
@@
$\mathcal{L}_{\mathrm{fkd}}=\beta_{1}\mathcal{L}_{\mathrm{fkd}_{1}}+\beta_{2}\mathcal{L}_{\mathrm{fkd}_{2}},$ (31)
@@
Finally, the total loss function combines the inner-layer feature distillation loss with the task loss $\mathcal{L}_{\mathrm{task}}$ and the soft label distillation loss $\mathcal{L}_{\mathrm{sdk}}$ as follows:
$\mathcal{L}=\gamma_{1}\mathcal{L}_{\mathrm{task}}+\gamma_{2}\mathcal{L}_{\mathrm{sdk}}+\gamma_{3}\mathcal{L}_{\mathrm{fkd}},$ (32)
@@ 4 Experiments
This chapter presents the implementation details, ablation studies, result comparisons, and visualizations. The details will be described in the following subsections.
@@ 4.1.1 Datasets
NTU-RGB+D [61] is a large-scale dataset comprising 60 action classes and 56,000 video clips. It includes detailed skeleton data covering various daily activities captured using a Kinect v2 sensor. The dataset provides 3D coordinates of 25 major body joints, enabling precise modeling of human movements. NTU-RGB+D is widely recognized as a standard benchmark in the action recognition field due to its variety of actions and large-scale data volume. The dataset is evaluated under two standard settings: Cross-Subject (XSub), where training and testing samples are divided based on the subjects performing the actions, and Cross-View (XView), where training and testing samples are divided based on the camera views.
NTU-RGB+D 120 [62], an extended version of NTU-RGB+D, significantly expands the dataset to include 120 action classes and over 114,000 video clips. This dataset offers richer skeleton data, making it one of the largest and most comprehensive resources for skeleton-based action recognition. Its inclusion of a broader range of actions and more extensive data volume provides an ideal foundation for developing and evaluating advanced action recognition algorithms. Similarly, this dataset is evaluated under two standard settings: Cross-Subject (XSub) and Cross-Setup (XSet), where training and testing samples are divided based on the camera setups used during data collection.
NW-UCLA [63] is a smaller dataset comprising 10 action classes and 1,494 video clips captured from multiple viewpoints. It provides 3D coordinates of 20 major body joints, offering diverse skeleton data for assessing the robustness of action recognition models to viewpoint variations. The dataset poses a challenge for models to perform consistently across different perspectives. The evaluation is conducted under the Cross-View setting, where training and testing samples are divided based on the camera angles used for data capture.
@@
Cross-Subject Cross-View (XSet)
@@
NTU RGB+D[61] 60 40,320 16,560 37,920 18,960 25
NTU RGB+D 120[62] 120 63,026 53,370 54,471 61,925 25
NW-UCLA[63] 10 - - 1,039 455 20
@@ 4.1.2 Implementation Details
The experiments are conducted on a server equipped with an NVIDIA SMI V100 GPU with 32 GB of memory. The models are implemented using PyTorch[65] and SpikingJelly[64], which provide efficient frameworks for spiking neural networks.
@@
The training configurations for MK-SGN across different datasets are summarized in Table 3. We adopt an initial learning rate of 0.1, a weight decay rate of 1e-4, and set the batch size based on the dataset: 64 for NTU-RGB+D, 64 for NTU-RGB+D 120, and 128 for NW-UCLA. Training is conducted for 64 epochs with a step decay schedule, reducing the learning rate to one-tenth after 50 epochs. In the testing phase, only the SGN model is evaluated. This ensures that the SGN’s performance reflects its learned capabilities without additional computational overhead.
@@ 4.2 Ablation Studies
To verify the effectiveness and efficiency of the proposed model, we first replace the activation layers of the classic one-stream 2S-AGCN model [21] with LIF neurons. This modified model, called Base-SGN*, is then trained and tested for comparison. Additionally, we compare our proposed model with Spikformer[36], a classic SNN-based transformer architecture. For a fair comparison, we adjust the layer structure and hidden dimension of Spikformer to match the MK-SGN architecture. These adjustments ensure that the comparison focuses on the modeling approach while minimizing the influence of structural differences. Based on these baselines, we conduct ablation and comparison experiments to evaluate the performance of MK-SGN.
@@ 4.2.1 Theoretical Energy Consumption Calculation
In the context of SNNs, synaptic operations ($\operatorname{SOPs}$) refer to the fundamental operations required to transmit spikes across synapses and perform computations within spiking neural networks. Specifically, $\operatorname{SOPs}$ are analogous to $\operatorname{FLOPs}$ in ANNs, serving as the core unit of the computational workload.
Each $\operatorname{SOP}$ corresponds to a spike-based accumulate ($\operatorname{AC}$) operation, which involves multiplying a binary spike (0 or 1) by a synaptic weight and accumulating the result in the post-synaptic neuron. This operation leverages the simplicity of binary spike-based computations, significantly reducing computational complexity compared to ANNs. In ANNs, operations are defined as Multiply-and-Accumulate ($\operatorname{MACs}$), where both multiplication and addition involve floating-point numbers. The use of $\operatorname{SOPs}$ in spiking neural networks thus highlights their efficiency, as $\operatorname{ACs}$ require fewer computational resources than $\operatorname{MACs}$.
As one of the critical components in our experimental evaluation, we calculate the SOPs for both Base-SGN* and MK-SGN to achieve two key objectives. First, this analysis underscores the advantages of our model on neuromorphic hardware by demonstrating its computational efficiency. Second, the calculated SOPs serve as the basis for estimating theoretical energy consumption, providing evidence of the energy-saving potential of MK-SGN.
@@
Method Architecture Param Spike Time Step SMF Knowledge Distillation SOPs FLOPs Power Ntu-RGB+D
(M) (G) (G) (mJ) Xsub(%) Xview(%)
ANN (10-layer)ST-GCN [20] 3.1 / / / / 3.48 16.01 81.5 88.3
@@
(6-layer)Spikformer[36]ICLR 2023 4.78 4 / / 1.69 24.07 2.17 73.9 80.6
(10-layer) MK-SGN 3.13 4 / / 0.96 10.43 0.86 72.9 79.6
(10-layer) MK-SGN 3.13 4 / 0.98 11.34 0.878 74.3 81.2
(10-layer) MK-SGN 3.24 4 / 1.00 11.45 0.912 74.2 80.1
(10-layer) MK-SGN 3.24 4 1.02 11.87 0.931 78.2 85.1
@@
Following [36, 37, 38], the $\operatorname{SOPs}$ for a given block or layer $l$ are defined as the product of the firing rate $r$, the number of spike time steps $S$, and the ($\operatorname{FLOPs}$) of the block or layer:
@@
$\quad+E_{\text{AC}}\times\bigg(\frac{k\times(k-1)}{2}\mathrm{SOP}_{\mathrm{SMIC}}+\sum_{n=n_{m}}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}$
@@
These formulations demonstrate how MK-SGN balances high performance with reduced theoretical energy consumption.
@@ 4.2.2 The Effect of Different Components
We use theoretical formulas to estimate the number of operations (OPs) required for classifying a single action sample to analyse the proposed MK-SGN’s energy efficiency and evaluate each component’s contribution within the model. Specifically, we calculate the FLOPs for ANN-based models, the FLOPs for SNN on traditional GPUs, and the theoretical SOPs on neuromorphic hardware. Based on these operation counts, we further compute the energy consumption per action sample during the inference phase.
Additionally, we evaluate the classification accuracy of MK-SGN through comprehensive ablation studies conducted on the NTU-RGB+D dataset. The results, including energy efficiency and accuracy comparisons, are summarized in Table 5.
@@
First, we validate the effectiveness of Base-SGN* as the basis for the ablation study. We compare it to the vanilla ST-GCN [20], calculating the theoretical number of operations (OPs) and energy consumption. The results show that Base-SGN* reduces the theoretical energy consumption to 0.536mJmJ per action sample. However, this increases the FLOPs and decreases the accuracy to $64.2$ and $71.3$. This indicates that simply replacing the activation layer is insufficient to maintain high classification accuracy.
@@
Building on Base-SGN*, we compare its performance with the proposed MK-SGN. Even without knowledge distillation, the single-modality MK-SGN achieves significant accuracy improvements of 10.3% on XSub and 9.6% on XView, with only a slight increase in OPs and theoretical power consumption. As shown in Table 5, the 10-layer MK-SGN achieves 72.9% accuracy on XSub and 79.6% on XView, with FLOPs at 10.43 G and power consumption at 0.86 mJ. Compared to ST-GCN, MK-SGN offers a better balance between accuracy and computational efficiency, significantly reducing energy consumption.
Next, we introduce the SMF module and knowledge distillation to further enhance performance. The 10-layer MK-SGN with SMF and knowledge distillation achieves 78.2% accuracy on XSub and 85.1% on XView, showcasing the benefits of SMF in improving accuracy while keeping SOPs at 1.02 G and power consumption at 0.931 mJ. Knowledge distillation further boosts accuracy with negligible impact on computational cost.
For the 6-layer MK-SGN, which eliminates redundant layers from the 10-layer architecture, the SMF and knowledge distillation configurations lead to an accuracy of 78.5% on XSub and 85.6% on XView. This version achieves low power consumption of 0.596 mJ and FLOPs of 7.76 G, demonstrating that the 6-layer model is the optimal design for balancing performance and efficiency.
The 4-layer configuration focuses on further reducing complexity by retaining only the most critical layers (such as channel upsampling layers). This 4-layer MK-SGN achieves lower accuracy—68.5% on XSub and 71.6% on XView—but still benefits from the SMF and knowledge distillation techniques, achieving improvements compared to the base model. However, the performance gain is not as substantial as in the 6-layer design, confirming that excessive layer reduction compromises the model’s representation capacity.
In summary, the addition of SMF and knowledge distillation enhances the accuracy of the model across all configurations. Among the tested architectures, the 6-layer MK-SGN strikes the best balance between accuracy and computational efficiency, demonstrating its suitability for real-time applications that require both high performance and low energy consumption.
@@
Table 7: Comparison of different fusion methods for MK-SGN. Accuracy and theoretical energy consumption are tested on NTU-RGB+D XSub.
Fusion Method Parameters Ops Power Accuracy
(M) (G) (mJ) (%)
SMF 2.17 0.67 0.596 76.1
Spiking Attention 2.24 0.71 0.613 76.0
Concatenation (cat) 1.90 0.64 0.589 75.4
Sum (direct addition) 1.89 0.63 0.579 75.2
  • 1.
    Note:cat is performed along the channel dimension with downsampling. Spiking Attention involves pairwise modality interaction followed by summation.
@@
In multi-modal classification, late fusion requires independent inference for each modality, causing energy consumption to increase by a factor of $n$ for $n$ modalities. For example, fusing four modalities—Joint (J), Bone (B), Joint Motion (JM), and Bone Motion (BM) in ST-GCN [20]—increases operations to $4\times 3.48\,\text{G}$ and power consumption to $4\times 16.01\,\text{mJ}$.
To reduce this energy cost while maintaining accuracy, we introduce the SMF module, which integrates modalities more efficiently. Table 6 shows the results from incremental fusion of modalities: J, J+B, J+B+JM, and J+B+JM+BM.
For the single Joint modality, the accuracy is $74.5\%$ with $0.573\,\text{mJ}$ of power consumption. As more modalities are added, accuracy increases to $76.1\%$ with a slight rise in energy consumption to $0.596\,\text{mJ}$. The number of parameters increases from $1.89\,\text{M}$ to $2.17\,\text{M}$, and operations from $0.64\,\text{G}$ to $0.67\,\text{G}$, showing that the SMF module efficiently integrates modalities without significantly increasing complexity.
These results confirm that the SMF module avoids the linear increase in energy consumption associated with late fusion, while improving accuracy by $1.6\%$ compared to single-modality fusion. This demonstrates the SMF module’s effectiveness in integrating multiple modalities efficiently.
Comparison with Other Fusion Methods
To further assess the effectiveness of the SMF module, we compare it with simpler fusion methods, including Spiking Attention, Concatenation (cat), and direct sum (Sum), using the same four modalities: Joint (J), Bone (B), Joint Motion (JM), and Bone Motion (BM). These methods are commonly used in multimodal fusion tasks but often suffer from higher computational overhead.
The comparison results are summarized in Table 7. Spiking Attention and Concatenation (cat) show similar accuracy improvements compared to single-modality fusion, with accuracies of $75.4\%$ and $76.0\%$, respectively. However, both methods require more computational resources than SMF. The Sum method, which directly adds the modality features, results in the lowest accuracy of $75.2\%$, as it fails to effectively model the interactions between modalities.
The SMF module achieves the highest accuracy of $76.1\%$ with relatively low increase in power consumption ($0.596\,\text{mJ}$), making it the most efficient method in terms of both accuracy and energy consumption. This validates the superiority of SMF over traditional late fusion techniques, as it balances accuracy improvements with computational efficiency.
@@
We evaluate the impact of our proposed knowledge distillation method by comparing SGN models trained with and without teacher guidance. The distillation method consists of Soft Label Distillation (SKD) and Inner-layer Feature Distillation (IFD). Ablation experiments are conducted to assess the contributions of each component. Additionally, we examine the effect of SKD using different modality combinations and the influence of IFD at various levels (low-level, high-level, or both).
The results in Table 8 demonstrate the effectiveness of SKD and IFD. With SKD, accuracy improves as more modalities are incorporated: single-modality achieves 75.1%, two-modalities 75.4%, and four-modalities 75.6%, highlighting the benefit of multimodal knowledge transfer. For IFD, focusing on high-level features yields better performance (75.3%) than low-level features (74.9%), indicating the significance of abstract information in guiding the student model.
Combining both SKD and IFD in the Full Knowledge Distillation (Full KD) setup achieves 75.9% accuracy, demonstrating the advantage of integrating soft-label guidance with feature-level distillation.
These results emphasize the importance of combining output-level and feature-level distillation to enhance SGN performance, with SKD aligning the student’s outputs with the teacher’s predictions, and IFD enabling the student to better emulate the teacher’s intermediate representations.
@@
Soft Label (Single Modal.) - J 75.1
@@
Table 9: Sensitivity Analysis of Model Hyperparameters and Accuracy for MK-SGN.
Model Hyperparameters Accuracy (%)
$\alpha_{b}$ $\alpha_{j}$ $\alpha_{bm}$ $\alpha_{jm}$ $\beta_{1}$ $\beta_{2}$ $\gamma_{1}$ $\gamma_{2}$ $\gamma_{3}$
0.1 0.1 0.1 0.1 0.5 0.5 0.5 0.5 0.5 77.1
0.5 0.5 0.5 0.5 0.5 0.5 1.0 1.0 1.0 77.5
0.25 0.5 0.25 0.25 0.5 0.5 0.8 0.9 1.0 78.3
0.75 0.25 0.75 0.75 0.5 0.5 0.6 0.7 1.0 77.9
0.25 0.25 0.5 0.5 0.5 0.5 0.9 1.0 1.0 78.0
0.25 0.25 0.25 0.25 0.5 0.5 1.0 1.0 1.0 78.5
0.25 0.25 0.25 0.25 1.0 1.0 0.7 0.6 0.5 77.6
0.1 0.25 0.25 0.25 1.0 0.5 0.8 0.9 0.6 77.5
@@
We carefully evaluate the contributions of each component in the MK-SGN model, with results summarized in Table 5. The fully integrated MK-SGN achieves accuracy of 78.5% and 85.6% on XSub and XView settings, respectively, with a spike time step of 4. Compared to Spikformer, MK-SGN reduces the parameter count by 54.6% and FLOPs by 67.7 %, significantly enhancing its feasibility for training on traditional GPUs. Moreover, MK-SGN decreases SOPs by 60.35% and theoretical energy consumption by 72.53% while achieving accuracy improvements of 4.6% and 5.0%, demonstrating its ability to balance high performance with low energy consumption.